OpenStudy (anonymous):

Find the radius of a cone with volume 199.23 cm3 and height 9 cm.

OpenStudy (anonymous):

use the formula in finding the volume of a cone

OpenStudy (anonymous):

The volume of a cone is given as\[(1/3)\pi*r^2*h\] . Use your 2 pieces of information and the formula to solve for r.

OpenStudy (anonymous):

no i am trying to find the radius not volume

OpenStudy (anonymous):

You are given the volume and the height in the question. If you plug those numbers in, you've got \[199.23=(1/3)\pi*r^2*9\] Do you see where 199.23 was plugged in for V and 9 was plugged into h (height)? Now you just need to solve for the only variable left in the equation, r, which is the radius.